# Recurrent Knowledge Identification and Fusion for Language Model Continual Learning

### v1

@@ Abstract
Continual learning (CL) is crucial for deploying large language models (LLMs) in dynamic real-world environments without costly retraining. While recent model ensemble and model merging methods guided by parameter importance have gained popularity, they often struggle to balance knowledge transfer and forgetting, mainly due to the reliance on static importance estimates during sequential training. In this paper, we present Recurrent-KIF, a novel CL framework for Recurrent Knowledge Identification and Fusion, which enables dynamic estimation of parameter importance distributions to enhance knowledge transfer. Inspired by human continual learning, Recurrent-KIF employs an inner loop that rapidly adapts to new tasks while identifying important parameters, coupled with an outer loop that globally manages the fusion of new and historical knowledge through redundant knowledge pruning and key knowledge merging. These inner-outer loops iteratively perform multiple rounds of fusion, allowing Recurrent-KIF to leverage intermediate training information and adaptively adjust fusion strategies based on evolving importance distributions. Extensive experiments on two CL benchmarks with various model sizes (from 770M to 13B) demonstrate that Recurrent-KIF effectively mitigates catastrophic forgetting and enhances knowledge transfer. [^1] The code will be released upon publication.
@@ 1 Introduction
Recently, model mixture-based methods have emerged as a mainstream approach for CL in LLMs Chen et al. (2023); Wu et al. (2024a); Rypeść et al. (2024). By leveraging parameter-efficient fine-tuning (PEFT) techniques, which reduce the computational burden, these methods can be broadly classified into two categories: model ensemble and model merging. Model ensemble methods assign a dedicated PEFT block to each task, capturing task-specific knowledge, which is then stored in a pool and dynamically selected during inference Zhu et al. (2024); Wang et al. (2024c). While effective, these methods require storing all task-specific models, leading to high memory consumption that grows with the number of tasks, which limits their scalability for long task sequences.
@@ 2.1 Continual Learning for LLMs
Model ensemble methods isolate parameters by assigning independent PEFT blocks to each task Feng et al. (2023); Pham et al. (2023); Ke et al. (2023); Li et al. (2024); He et al. (2024); Wang et al. (2024a). For example, O-LoRA Wang et al. (2023a) enforces orthogonality among LoRA adapters, while SAPT Zhao et al. (2024) uses a selection module to combine blocks based on task correlations. While preserving task-specific knowledge, they hinder inter-task transfer and incur high memory overhead as the number of tasks increases, limiting their scalability.
In contrast, model merging methods combine multiple models into a single model Cheng et al. (2024); Alexandrov et al. (2024); Ren et al. (2024), alleviating memory constraints. For example, global model merging approaches Wortsman et al. (2022); Ilharco et al. (2023) perform a weighted fusion of models before and after training, typically assuming that all model weights contribute equally to each task. However, determining which and how to merge parameters remains an open problem. In this paper, we propose Recurrent-KIF, a novel framework that leverages the dynamic importance of parameters across different tasks by employing knowledge identification and fusion techniques to mitigate CF and promote KT.
@@ 2.2 Parameter Importance Identification
Identifying important parameters or knowledge regions within LLMs has gained significant attention in the NLP community Zhao et al. (2023); Liu et al. (2023); Feng et al. (2024a); Xu et al. (2024); Shi et al. (2024). This research improves our understanding of LLMs and enhances their performance across a variety of tasks, including model editing Wang et al. (2024d), compression Zhang et al. (2023).
@@ 3.2 Outer Learner with Knowledge Fusion
The outer loop manages the global merging of knowledge, guided by parameter importance. To access historical knowledge, after acquiring $\theta_{b(Q)}$, the outer loop samples data $\phi_{b}$ from the memory buffer $\mathcal{M}_{<k}$. It then performs a single training iteration, updating the parameters to $\theta_{b(M)}$. Then the outer task vector $\tau_{b}^{\text{out}}\in\mathbb{R}^{n}$, capturing historical task information, is defined as:
@@ 4 Experiments and Analysis
We evaluate Recurrent-KIF using two distinct language model architectures: the encoder-decoder T5 model Raffel et al. (2020) (T5-large and T5-xl), and the decoder-only LLaMA model Touvron et al. (2023) (LLaMA2-7B and LLaMA2-13B) [^3] Due to being limited by an academic computing budget, we employed 8-bit quantization for LLaMA models.. Hyperparameters $\alpha_{1}$ and $\alpha_{2}$ in Eq. (4) and Eq. (7) are set to 0.55, with the number of inner loop iterations $Q$ set to 8. Following Zhao et al. (2024), 2% of the original training set is used for replay samples. All experiments are averaged over 3 runs. More details are in Appendix D.
@@ 4.1 Main Results
Moreover, when compared to parameter importance-based methods like TaSL and VR-MCL, Recurrent-KIF consistently delivers the best OP and BWT scores. Notably, Recurrent-KIF outperforms the state-of-the-art CL method, MIGU, increasing OP from 76.6% to 78.1%. These results underscore the effectiveness of our recurrent knowledge identification and fusion framework, validating the advantages of dynamic parameter importance estimation in mitigating CF and promoting KT.
@@ Appendix C Dataset Statistics
Table 11 show details of the datasets we used for our experiments, along with their evaluation metrics Wang et al. (2023a); Feng et al. (2024c); Xu et al. (2023). For the Long Sequence benchmark, this includes five tasks from the standard CL benchmark (AG News, Amazon reviews, Yelp reviews, DBpedia and Yahoo Answers), four from GLUE benchmark (MNLI, QQP, RTE, SST2), five from SuperGLUE benchmark (WiC, CB, COPA, MultiRC, BoolQ), and the IMDB movie reviews dataset. We report 6 different task orders used for our experiments in Table 12. Table 13 shows prompts for different tasks. NLI denotes natural language inference Lu et al. (2021b), including MNLI, RTE and CB. SC denotes sentiment analysis, including Amazon, Yelp, SST-2 and IMDB. TC denotes topic classification, including AG News, Dbpedia and Yahoo.
@@ Appendix E Algorithm
0:  Current task dataset $\mathcal{D}_{k}$, memory buffer $\mathcal{M}_{<k}$, model weights $\theta$, initial inner-loop learning rate $\beta_{in}$, outer-loop learning rate $\beta_{out}$, number of inner-loop steps $Q$, hyperparameters $\alpha_{1},\alpha_{2}$, total training step $N$.
@@
15:     sample memory data $\phi_{b}$ from $\mathcal{M}_{<k}$
16:     $\theta_{b(M)}=\theta_{b(Q)}-\beta_{out}\nabla\mathcal{L}\left(\theta_{b(Q)}\right)$
17:     calculate $\tau_{b}^{\text{out}}=\theta_{b(M)}-\theta_{b(Q)}$
18:     calculate $I_{b}^{\text{out}}$ via Eq. (7)
19:     # knowledge fusion.
20:     obtain $m_{b}^{\text{in}}$, $m_{b}^{\text{out}}$ via Eq. (8)
21:     update $\theta_{b+1}=\theta_{b}+(m_{b}^{\text{in}}\odot\tau_{b}^{\text{in}}+m_{b}^{\text{out}}\odot\tau_{b}^{\text{out}})$
22:  end for

### v2

@@ Abstract
Continual learning (CL) is crucial for deploying large language models (LLMs) in dynamic real-world environments without costly retraining. While recent model ensemble and model merging methods guided by parameter importance have gained popularity, they often struggle to balance knowledge transfer and forgetting, mainly due to the reliance on static importance estimates during sequential training. In this paper, we present Recurrent-KIF, a novel CL framework for Recurrent Knowledge Identification and Fusion, which enables dynamic estimation of parameter importance distributions to enhance knowledge transfer. Inspired by human continual learning, Recurrent-KIF employs an inner loop that rapidly adapts to new tasks while identifying important parameters, coupled with an outer loop that globally manages the fusion of new and historical knowledge through redundant knowledge pruning and key knowledge merging. These inner-outer loops iteratively perform multiple rounds of fusion, allowing Recurrent-KIF to leverage intermediate training information and adaptively adjust fusion strategies based on evolving importance distributions. Extensive experiments on two CL benchmarks with various model sizes (from 770M to 13B) demonstrate that Recurrent-KIF effectively mitigates catastrophic forgetting and enhances knowledge transfer. [^1] https://github.com/WoodScene/Recurrent_KIF
@@ 1 Introduction
Recently, model mixture-based methods have emerged as a mainstream approach for CL in LLMs Chen et al. (2023); Wu et al. (2024a); Rypeść et al. (2024); Chen et al. (2024). By leveraging parameter-efficient fine-tuning (PEFT) techniques, which reduce the computational burden, these methods can be broadly classified into two categories: model ensemble and model merging. Model ensemble methods assign a dedicated PEFT block to each task, capturing task-specific knowledge, which is then stored in a pool and dynamically selected during inference Shengyuan et al. (2023); Zhu et al. (2024); Wang et al. (2024c). While effective, these methods require storing all task-specific models, leading to high memory consumption that grows with the number of tasks, which limits their scalability for long task sequences.
@@ 2.1 Continual Learning for LLMs
Model ensemble methods isolate parameters by assigning independent PEFT blocks to each task Feng et al. (2023); Pham et al. (2023); Ke et al. (2023); Li et al. (2024); He et al. (2024); Wang et al. (2024a); Zhang et al. (2025a). For example, O-LoRA Wang et al. (2023a) enforces orthogonality among LoRA adapters, while SAPT Zhao et al. (2024) uses a selection module to combine blocks based on task correlations. While preserving task-specific knowledge, they hinder inter-task transfer and incur high memory overhead as the number of tasks increases, limiting their scalability.
In contrast, model merging methods combine multiple models into a single model Cheng et al. (2024); Alexandrov et al. (2024); Ren et al. (2024); Shengyuan et al. (2023); Zhang et al. (2025a), alleviating memory constraints. For example, global model merging approaches Wortsman et al. (2022); Ilharco et al. (2023) perform a weighted fusion of models before and after training, typically assuming that all model weights contribute equally to each task. However, determining which and how to merge parameters remains an open problem. In this paper, we propose Recurrent-KIF, a novel framework that leverages the dynamic importance of parameters across different tasks by employing knowledge identification and fusion techniques to mitigate CF and promote KT.
@@ 2.2 Parameter Importance Identification
Identifying important parameters or knowledge regions within LLMs has gained significant attention in the NLP community Zhao et al. (2023); Liu et al. (2023); Feng et al. (2024a); Xu et al. (2024); Shi et al. (2024); Zhang et al. (2025b). This research improves our understanding of LLMs and enhances their performance across a variety of tasks, including model editing Wang et al. (2024d); Feng et al. (2025), compression Zhang et al. (2023); Jiang et al. (2023).
@@ 3.2 Outer Learner with Knowledge Fusion
The outer loop manages the global merging of knowledge, guided by parameter importance. To access historical knowledge, after acquiring $\theta_{b(Q)}$, the outer loop samples data $\phi_{b}$ from the memory buffer $\mathcal{M}_{<k}$. It then performs several training iterations, updating the parameters to $\theta_{b(M)}$. Then the outer task vector $\tau_{b}^{\text{out}}\in\mathbb{R}^{n}$, capturing historical task information, is defined as:
@@ 4 Experiments and Analysis
We evaluate Recurrent-KIF using two distinct language model architectures: the encoder-decoder T5 model Raffel et al. (2020) (T5-large and T5-xl), and the decoder-only LLaMA model Touvron et al. (2023) (LLaMA2-7B and LLaMA2-13B) [^3] Due to being limited by an academic computing budget, we employed 8-bit quantization for LLaMA models.. Hyperparameters $\alpha_{1}$ and $\alpha_{2}$ in Eq. (4) and Eq. (7) are set to 0.55, with the number of inner loop iterations $Q$ set to 8, and the number of outer loop iterations set to 4. Following Zhao et al. (2024), 2% of the original training set is used for replay samples. All experiments are averaged over 3 runs. More details are in Appendix D.
@@ 4.1 Main Results
Moreover, when compared to parameter importance-based methods like TaSL and VR-MCL, Recurrent-KIF consistently delivers the best OP and BWT scores. Notably, Recurrent-KIF outperforms the state-of-the-art CL method, MIGU, increasing OP from 76.6% to 78.1%. These results underscore the effectiveness of our recurrent knowledge identification and fusion framework, validating the advantages of dynamic parameter importance estimation in mitigating forgetting and promoting knowledge transfer.
@@ 5 Conclusion
Acknowledgements
We thank the anonymous reviewers for their valuable feedback and the support of the PolyU Research Student Attachment Programme.
@@ Appendix C Dataset Statistics
Table 11 show details of the datasets we used for our experiments, along with their evaluation metrics Wang et al. (2023a); Feng et al. (2024c); Xu et al. (2023). For the Long Sequence benchmark, this includes five tasks from the standard CL benchmark (AG News, Amazon reviews, Yelp reviews, DBpedia and Yahoo Answers), four from GLUE benchmark (MNLI, QQP, RTE, SST2), five from SuperGLUE benchmark (WiC, CB, COPA, MultiRC, BoolQ), and the IMDB movie reviews dataset. We report 6 different task orders used for our experiments in Table 12. Table 13 shows prompts for different tasks. NLI denotes natural language inference Lu et al. (2021b), including MNLI, RTE and CB. SC denotes sentiment analysis, including Amazon, Yelp, SST-2 and IMDB. TC denotes topic classification, including AG News, Dbpedia and Yahoo Xu et al. (2025); Chen et al. (2024).
@@ Appendix E Algorithm
0:  Current task dataset $\mathcal{D}_{k}$, memory buffer $\mathcal{M}_{<k}$, model weights $\theta$, initial inner-loop learning rate $\beta_{in}$, outer-loop learning rate $\beta_{out}$, number of inner-loop steps $Q$, number of outer-loop steps $S$, hyperparameters $\alpha_{1},\alpha_{2}$, total number of fusion steps $N$.
@@
15:     $\theta_{b(M)}^{0}=\theta_{b(Q)}$
16:     for $s=1$ to $S$ do
17:        sample memory data $\phi_{b}^{s}$ from $\mathcal{M}_{<k}$
18:        $\theta_{b(M)}^{s}=\theta_{b(M)}^{s-1}-\beta_{out}\nabla\mathcal{L}\left(\theta_{b(M)}^{s-1}\right)$
19:     end for
20:     calculate $\tau_{b}^{\text{out}}=\theta_{b(M)}^{S}-\theta_{b(Q)}$
21:     calculate $I_{b}^{\text{out}}$ via Eq. (7)
22:     # knowledge fusion.
23:     obtain $m_{b}^{\text{in}}$, $m_{b}^{\text{out}}$ via Eq. (8)
24:     update $\theta_{b+1}=\theta_{b}+(m_{b}^{\text{in}}\odot\tau_{b}^{\text{in}}+m_{b}^{\text{out}}\odot\tau_{b}^{\text{out}})$
25:  end for